How do you find the LCM of a set of numbers 16 AND 20
Answer by stanbon(75887)

LCM of a set of numbers 16 AND 20
----
List the prime-factor form of each:
16 = 2^4
20 = 2^2*5
----
The LCM has each of the prime factors is them highest power.
---
LCM = 2^4*5 = 16*5 = 80
=========================
Cheers,
Stan H.
===================